### v3

@@ Abstract
Deep reinforcement learning (RL) policies, although optimal in terms of task rewards, may not align with the personal preferences of human users. To ensure this alignment, a naïve solution would be to retrain the agent using a reward function that encodes the user’s specific preferences. However, such a reward function is typically not readily available, and as such, retraining the agent from scratch can be prohibitively expensive. We propose a more practical approach - modify the already trained policy to accommodate the user-specific needs with the help of human feedback. To this end, we infer the user’s intent through trajectory-level feedback and combine it with the trained task policy via a theoretically grounded dynamic policy fusion approach. As our approach collects human feedback on the very same trajectories used to learn the task policy, it does not require any additional interactions with the environment, making it a zero-shot approach in terms of environment interactions. We empirically demonstrate in a number of environments that our proposed dynamic policy fusion approach consistently achieves the intended task while simultaneously adhering to user-specific needs.
@@ 1 Introduction
Reinforcement learning (RL) has demonstrated its effectiveness in various real-world scenarios, including computer games [16], inventory management, autonomous driving, robotics [8], healthcare [31], recommendation systems [26] etc. In RL, the learning agent typically learns to maximise its reward by learning an optimal task policy through interactions with the environment. However, in certain scenarios, a user may desire a policy that subtly deviates from this optimal policy to accommodate their personal preference or style. For example, in a navigation setting, users may prefer to take quieter, less congested routes, even if this means slightly longer travel times compared to the route determined to be the optimal one by a reinforcement learning (RL) agent trained to minimize travel time. This preference allows users to enjoy a more relaxed and stress-free journey.
A simple solution to accommodate such preferences in a policy would be to retrain an RL agent from scratch, using a reward function that takes users’ preferences into account. However, such a reward function could be challenging to obtain. In addition, the poor sample efficiency of RL could render this approach infeasible, as it would require the agent to interact with the environment again, which could be expensive in terms of time, energy, etc., We posit that a more practical approach would be to adapt the originally trained policy by personalising it based on a user’s specific preferences on the trajectories already collected during training. This approach to personalisation eliminates the need for additional agent-environment interactions, thereby making it a practically feasible option for adapting already trained policies.
In this work, we present a policy fusion-based approach to learn such a Personalised policy that respects human preferences while also satisfying the task objective. To infer human preferences, the trajectories used to initially learn the task (we refer to the corresponding policy as the Task-specific policy) are numerically scored to reflect the degree to which they align with the human’s preferred behaviour. These trajectory-score pairs are fed to an LSTM [10] to construct an Intent-specific policy. We choose an LSTM-based approach as it is well-suited to model temporal patterns within the trajectories. Once the user’s intent is inferred, the key idea is then to fuse the task-specific policy with this intent-specific policy, such that the resulting personalised policy completes the task at hand while simultaneously respecting user-specific needs. We design the described policy fusion method to theoretically ensure that the divergence of the personalised policy with respect to the task-specific policy remains bounded implying that the personalised policy upon fusion cannot change arbitrarily. Since the inference of the user’s intent (learning the intent-specific policy) is done by reusing the same trajectories used for training the task policy/task-specific policy, our approach obviates the need for any additional environment interactions, making it a zero-shot scheme.
Despite the mentioned advantages of the described static policy fusion-based approach, we show that this policy fusion approach is insufficient, as it does not account for whether one policy is over/under-represented, which can lead to undesirable effects. For example, in a navigation scenario, the preference for visiting a particular state may cause the agent to repeatedly visit that state, thereby ignoring the primary navigation objective. In other words, static policy fusion may cause one of the policies to dominate the other, leading to unwanted agent behaviours. We address this problem through a novel dynamic policy fusion approach that automatically regulates the dominance of the intent-specific policy by regulating the associated temperature parameter of the Boltzmann distribution [23] used to represent the policy. This helps control the relative contribution of the intent-specific policy in the personalised policy, thereby allowing the agent to complete the task while respecting user-specific needs/preferences.
We note that the problem of respecting additional user-specific objectives could also be formulated as a multi-objective reinforcement learning [22, 9, 15] problem, although such approaches are applicable only when rewards corresponding to each objective are available during learning. As such, the present work aims to deal with how one can adopt a trained policy after one has learned an initial task.
Some of our early work in this direction has been reported in [25] and our current work provides a comprehensive presentation including improvements in the method, new theoretical analysis to firmly ground the main intuition behind the policy fusion method, and extensive evaluation across diverse environments.
@@ 2 Related Work
A number of works in interactive reinforcement learning [29, 14] examine how agents can learn from state-action level human feedback. However, such methods are limited, as providing state-action-level feedback is non-intuitive and cognitively taxing. Preference-based reinforcement learning [5, 12], uses trajectory-level human preference feedback, using which a corresponding reward function is learned. However, following the learning of the reward function, this approach may require the agent to interact with the environment again to learn the corresponding policy. As our present work focuses on developing a zero-shot approach, to estimate human intent, we instead use RUDDER [1], an LSTM-based approach for credit assignment by directly approximating the Q-values from agent trajectories. These Q-values are then translated into the required intent-specific policy. There are more recent work addressing credit assignment such as [19, 32], but we resort to RUDDER for our work due to it’s simplicity.
@@
In the context of policy fusion, a closely related work by Sestini et al. [24] combines policies with different fusion methods. However, the described policy fusion methods are static and may result in the dominance of one of the constituent policies. Policy fusion has been explored previously in Haarnoja et al. [7] and Hunt et al. [11]. Such works study the concept of learning different policies independently using soft Q-functions, each with its own reward function. These individual policies are later combined, leading to the emergence of new behaviours in robotic manipulations. Our work views policy fusion through a lens of personalisation, requiring the fusion to satisfy certain constraints. In addition, unlike existing policy fusion methods, our dynamic policy fusion approach prevents the over-dominance of the component policies.
@@
The overarching aim of our work is to personalise existing task policies using trajectory level human feedback without the need for additional environment interactions. The goal of personalisation could possibly be viewed from a multi-objective reinforcement learning (MORL) [9] or constrained reinforcement learning [28] perspective. Multi-objective reinforcement learning (MORL) embodies personalisation by determining which objectives take precedence, and addressing conflicting objectives [2] through Pareto-optimal solutions. In Mo et al. [17], a personalised dialogue system is developed using transfer learning, adapting common dialogue knowledge from a source domain to a target user, whereas Bodas et al. [4] aims to personalise non-player characters (NPCs) to human skill levels to enhance player engagement. This approach designs a composite scalar reward function that implicitly matches the NPC’s skill level to that of the human player. However, it necessitates additional environment interaction and NPC training with the engineered reward function. Unlike these works, our approach does not require any additional interaction or specifically crafting the reward function. We believe these benefits make the present work particularly relevant for real-world scenarios.
@@ 3 Preliminaries and Background
We briefly survey some of the related backgrounds that form the basis for our work. We refer to the policy trained to solve the task as the task-specific policy $(\pi_{\phi})$ and the policy that captures human intent as the intent-specific policy$(\pi_{\psi})$.
@@ 3.1 Reinforcement Learning
The policies considered in this work are learned by solving tasks represented as individual MDPs [21] $\mathcal{M}=(\mathcal{S},\mathcal{A},P,R,\gamma)$ where $\mathcal{S}$, and $\mathcal{A}$ are the state and action spaces. $P$ is the transition function that captures the transition dynamics of the environment. $R$ is the reward function and $\gamma$ is the discount factor. In each task, $R$ can vary, while the remaining components stay the same. At timestep $t$, the agent in state $s_{t}\in\mathcal{S}$ takes an action $a_{t}\in\mathcal{A}$ and obtains a reward $R(s_{t+1},s_{t},a_{t})$ and moves to state $s_{t+1}$ according to the transition function $P(s_{t+1}|s_{t},a_{t})$. A policy $\pi(a_{t}|s_{t})$ outputs the probability of taking an action $a_{t}$ from a given state $s_{t}$. The episodic discounted return is $G_{t}=\sum_{t=0}^{T}\gamma^{t}r_{t}$, where $\gamma$ specifies how much the future reward is discounted and $T$ is the total number of timesteps. The agent’s objective is to maximise the future expected reward $\mathbb{E}[G_{t}]$ by learning a Q-function $Q(s_{t},a_{t})$, which estimates the expected cumulative reward.
@@ 4 Methodology
We assume that the trajectory data used to train task-specific policy is accessible, and reusing this data allows us to perform zero-shot personalisation i.e., without collecting new environment interactions. A subset of this data is sampled with personalised human feedback scores with more desirable trajectories assigned higher scores. We then infer the human intent (intent-specific policy) using RUDDER (Section 3.2) and dynamically fuse it with the task-specific policy. The overview of our method is illustrated in Figure 1. The subsequent sections provide in-depth details of the components of our personalisation approach.
@@ 4.1 Learning Human Intent using LSTM
At each time step, the state-action vector is fed into the LSTM units. In the case of image inputs, we pre-train a Variational Autoencoder (VAE) to reduce the dimensionality of the state. We use FiLM [20] to modulate the state vector with the action vector. In [20], a FiLM network conditions the feature map of the neural network depending on another input signal. Here, we condition the state feature with the action vector and train the LSTM using the human feedback data as training labels. With this training setup, (More details in supplementary Section B and C) the model approximates the Q-value, which is then converted to the intent-specific policy as described in the next section.
@@ 4.2 Policy Construction
We use a DQN parameterised by $\phi$ to learn the task and it produces the Q-values ($Q$) for each state-action pair. Similarly, the LSTM parameterised by $\psi$ also produces the Q-values ($Q^{\prime}$) corresponding to the human intent. We choose the Boltzmann distribution to convert the Q-values into a policy. Therefore $\pi_{\phi}$ is the task-specific policy and $\pi_{\psi}$ is the intent-specific policy. This choice is motivated by the fact that if the intersection of the support of two policies is an empty set, their product would result in a random policy. By contrast, the Boltzmann distribution assigns at least a small probability to all actions, ensuring that the support of the distribution covers the entire action space.
@@ 4.3 Policy Fusion
Policy fusion is a process of combining two or more policies to produce a new policy. Sestini et al. [24] discussed several approaches through which policy fusion can be performed. We also aim to obtain a personalised policy by fusing the intent-specific and task-specific policies. However, policy fusion in the context of personalisation should satisfy two constraints. Firstly, fused policy should be identical to the task-specific policy if the task-specific and intent-specific policies are identical to each other - a property that we refer to as the invariability constraint. This constraint is crucial because it ensures that when the objectives of the task and the human are aligned, the personalised policy remains consistent and does not deviate during the fusion of identical constituent policies. Correspondingly, we define invariant policies as follows:
@@
Two policies $\pi_{1}(a|s)$ and $\pi_{2}(a|s)$ are said to be invariant policies if $KL\left(\pi_{1}(a|s)\middle\|\pi_{2}(a|s)\right)=0\ \forall\ s\in\mathcal{S}$.
@@
If $\pi_{\phi}(a|s)=\pi_{\psi}(a|s)$ then $KL\left(\pi_{\phi}(a|s)\middle\|\pi_{f}(a|s)\right)=KL\left(\pi_{\phi}(a|s)\middle\|\pi_{\phi}(a|s)\right)=0$
@@
$+\dfrac{1}{2}\log\left(\zeta\right)$
The above theorem establishes the boundedness of the divergence between the task-specific policy and the personalised policy. Nevertheless, we can consider other choices for policy fusion as studied in the previous work [24], such as the weighted average or the product of the two policies. However, the former does not operate on the common support, and the latter does not satisfy the invariability constraint, even when the policies are exactly the same, as evidenced in the Lemma 2. For a detailed analysis, we refer the reader to the supplementary Section E.
@@
Let $\pi_{\psi}$ and $\pi_{\phi}$ be intent-specific policy and task-specific policies respectively and $\pi_{\psi}$ is not a random policy. Then $\pi_{\phi}$ and $\tfrac{1}{Z}\pi_{\psi}\pi_{\phi}$, where $Z$ is normalising factor are not invariant policies in any condition.
@@
$KL\left(\pi_{\phi}|\tfrac{\pi_{\phi}\pi_{\psi}}{Z}\right)=$ $\sum_{a}\pi_{\phi}\log\left(\dfrac{\pi_{\phi}Z}{\pi_{\phi}\pi_{\psi}}\right)$
$=$ $\sum_{a}\pi_{\phi}\log\left(\dfrac{Z}{\pi_{\psi}}\right)$
$=$ $\log\left(Z\right)-\sum_{a}\pi_{\phi}\log\pi_{\psi}$
Where $Z$ is the normalising factor, $Z=\sum_{a}\pi_{\phi}\pi_{\psi}$. Assume the expression above is 00 $\Rightarrow\log\left(Z\right)=\sum_{a}\pi_{\phi}\log\pi_{\psi}$
$\log\left(\sum_{a}\pi_{\phi}\pi_{\psi}\right)=\sum_{a}\pi_{\phi}\log\pi_{\psi}$
$\Rightarrow\log\left(E_{\pi_{\phi}}\left(\pi_{\psi}\right)\right)=E_{\pi_{\phi}}\left(\log\left(\pi_{\psi}\right)\right)$
The Jensen’s Inequality states that $E_{\pi_{\phi}}\left(\log\left(\pi_{\psi}\right)\right)\leq\log\left(E_{\pi_{\phi}}\left(\pi_{\psi}\right)\right)$ and this expression is an equality only if $\pi_{\psi}$ is a random policy, which is a contradiction. ∎
@@ 4.4 Pitfalls of Static Fusion
1:while not $done$ do
2:     $Q_{t}\leftarrow[]$triangleright DQN Q-values reset the vector to null
3:     $Q^{\prime}_{t}\leftarrow[]$triangleright LSTM Q-values reset the vector to null
4:     $r_{t}\leftarrow[]$triangleright Human-induced reward reset the vector to null
5:     $Q_{t}\leftarrow\phi(s_{t})$triangleright Invoking DQN
6:     $Q^{\prime}_{t}\leftarrow\phi(s_{t},a)$triangleright Invoking LSTM for each action
@@
8:     $\pi_{\phi}\leftarrow BoltzmanDistribution(Q_{t},T_{\phi})$triangleright Task-specific policy
9:     $\pi_{\psi}\leftarrow BoltzmanDistribution(Q^{\prime}_{t},T_{\psi})$triangleright Intent-specific policy
10:     $a_{t}\leftarrow\arg\max_{a\in\mathcal{A}}\sqrt{\pi_{\phi}(a|s_{t})\times\pi_{\psi}(a|s_{t})}$triangleright Policy fusion
11:     $r^{\prime}\leftarrow r_{t}-mean(r_{t})$triangleright Adjusting as in Eq (5)
12:     $g(t)\leftarrow\sum_{t^{\prime}=0}^{t}r^{\prime}(s_{t^{\prime}},a_{t^{\prime}})$triangleright Accumulated shifted human-induced reward
13:     $T\leftarrow max(T_{min},\frac{T_{max}}{1+exp(-m(g-\eta))})$triangleright Updating $T$
@@
15:     $t\leftarrow t+1$triangleright Increment the time-step

### v4

@@ Abstract
Deep Reinforcement Learning policies, although optimal in terms of task rewards, may not align with the personal preferences of human users. To ensure this alignment, a naïve solution would be to retrain the agent using a reward function that encodes the user’s specific preferences. However, such a reward function is typically not readily available, and as such, retraining the agent from scratch can be prohibitively expensive. We propose a more practical approach - modify the already trained policy to accommodate the user-specific needs with the help of human feedback. To this end, we infer the user’s intent through trajectory-level feedback and combine it with the trained task policy via a theoretically grounded dynamic policy fusion approach. Our approach collects human feedback on the same trajectories used to learn the task policy. As a result, it does not require any additional interactions with the environment. This makes it a zero-shot approach regarding environment interactions. We empirically demonstrate in a number of environments that our proposed dynamic policy fusion approach consistently achieves the intended task while simultaneously adhering to user-specific needs.
@@ 1 Introduction
Reinforcement learning (RL) has achieved remarkable success across a variety of domains, including games [14], robotics [8], healthcare [27], and recommendation systems [22]. Standard RL agents are optimised to maximise a predefined reward signal, which typically encodes task efficiency. While this approach ensures task completion, the resulting agent behaviour may not align with the subjective preferences of human users. For instance, in a driving context, Reinforcement Learning (RL) agent trained to minimise travel time might adopt aggressive manoeuvres, whereas a human passenger would likely prefer a smoother, safer driving style. Since such personal preferences are rarely captured in the original reward function, aligning agent behaviour with user expectations remains a critical challenge.
A common approach involves retraining the agent with a reward function modified to incorporate user preferences. However, this is often impractical; designing such personalised reward functions is non-trivial, and retraining from scratch demands substantial additional interaction with the environment, which can be costly or infeasible. A more pragmatic alternative is to adapt a pre-trained policy which we refer to as the task-specific policy so that it continues to achieve the primary task goal while also accommodating the user’s intent.
Our framework personalises a pre-trained task-specific policy by inferring user intent from trajectory-level feedback. Critically, this feedback is applied to the same trajectories generated during the training of the task policy itself, enabling a zero-shot approach that requires no new environment interaction. From this feedback, we construct an auxiliary intent-specific policy that captures user preferences on how the task should be performed. The final, personalised policy is formed by fusing the original task-specific policy with this new intent-specific policy. A key challenge in this process is to balance the two objectives: blindly following user intent could compromise task success, while ignoring it defeats the purpose of personalisation. To resolve this, we introduce dynamic policy fusion mechanism that adaptively modulates the influence of each policy based on the current state, ensuring both task completion and alignment with user preferences.
The main contributions of this paper are as follows:
  • We propose a zero-shot personalisation framework that adapts pre-trained RL policies to user-specific preferences using trajectory-level feedback, eliminating the need for additional environment interaction.
  • We introduce a policy fusion formulation with formal guarantees, including bounded divergence from the original task policy and an invariability property for identical constituent policies.
  • We develop a mechanism to dynamically adapt the influence of the intent policy, ensuring the final behaviour robustly balances task completion with user alignment.
This paper is an extension of our preliminary work presented in [21]. We expand upon the initial research in three key aspects:
  • We provide a comprehensive theoretical analysis of the policy fusion mechanism, establishing formal guarantees for its behaviour.
  • Establishing the necessity of dynamic policy fusion through detailed empirical studies.
  • We broaden the experimental evaluation to include more diverse human preference models, validating the robustness and versatility of our framework.
The rest of this paper is organised as follows: Section 2 reviews related work and Section 3 details the necessary preliminaries. Section 4 presents our proposed approach, followed by our experimental evaluations in Section 5. Finally, Section 6 discusses the limitations of our work and Section 7 concludes the paper.
@@ 2 Related Work
A number of works in interactive RL [25, 13] examine how agents can learn from state-action level human feedback. However, such methods are limited, as providing state-action-level feedback is non-intuitive and cognitively taxing. Preference-based RL [5, 11], uses trajectory-level human preference feedback, using which a corresponding reward function is learned. However, following the learning of the reward function, this approach may require the agent to interact with the environment again to learn the corresponding policy. As our present work focuses on developing a zero-shot approach, to estimate human intent, we instead use Return Decomposition for Delayed Rewards (RUDDER) [1], an Long Short Term Memory (LSTM)-based approach for credit assignment by directly approximating the Q-values from agent trajectories. These Q-values are then translated into the required intent-specific policy. There are more recent work addressing credit assignment such as [17, 28], but we resort to RUDDER for our work due to it’s simplicity.
@@
In the context of policy fusion, a closely related work by [20] combines policies with different fusion methods. However, the described policy fusion methods are static and may result in the dominance of one of the constituent policies. Policy fusion has been explored previously in [7] and [10]. Such works study the concept of learning different policies independently using soft Q-functions, each with its own reward function. These individual policies are later combined, leading to the emergence of new behaviours in robotic manipulations. Our work views policy fusion through a lens of personalisation, requiring the fusion to satisfy certain constraints. In addition, unlike existing policy fusion methods, our dynamic policy fusion approach prevents the over-dominance of the component policies.
@@
The overarching aim of our work is to personalise existing task policies using trajectory level human feedback without the need for additional environment interactions. The goal of personalisation could possibly be viewed from a multi-objective reinforcement learning (MORL) [9] or constrained RL [24] perspective. MORL embodies personalisation by determining which objectives take precedence, and addressing conflicting objectives [2] through Pareto-optimal solutions. In  [15], a personalised dialogue system is developed using transfer learning, adapting common dialogue knowledge from a source domain to a target user, whereas  [4] aims to personalise non-player characters (NPCs) to human skill levels to enhance player engagement. This approach designs a composite scalar reward function that implicitly matches the NPC’s skill level to that of the human player. However, it necessitates additional environment interaction and NPC training with the engineered reward function. Unlike these works, our approach does not require any additional interaction or specifically crafting the reward function. We believe these benefits make the present work particularly relevant for real-world scenarios.
@@ 3 Preliminaries and Background
[…]
@@ 3.1 Reinforcement Learning
The policies considered in this work are learned by solving tasks represented as individual Markov Decision Processes (MDPs) [19] $\mathcal{M}=(\mathcal{S},\mathcal{A},P,R,\gamma)$ where $\mathcal{S}$, and $\mathcal{A}$ are the state and action spaces. $P$ is the transition function that captures the transition dynamics of the environment. $R$ is the reward function and $\gamma$ is the discount factor. In each task, $R$ can vary, while the remaining components stay the same. At timestep $t$, the agent in state $s_{t}\in\mathcal{S}$ takes an action $a_{t}\in\mathcal{A}$ and obtains a reward $R(s_{t+1},s_{t},a_{t})$ and moves to state $s_{t+1}$ according to the transition function $P(s_{t+1}|s_{t},a_{t})$. A policy $\pi(a_{t}|s_{t})$ outputs the probability of taking an action $a_{t}$ from a given state $s_{t}$. The episodic discounted return is $G_{t}=\sum_{t=0}^{T}\gamma^{t}r_{t}$, where $\gamma$ specifies how much the future reward is discounted and $T$ is the total number of timesteps. The agent’s objective is to maximise the future expected reward $\mathbb{E}[G_{t}]$ by learning a Q-function $Q(s_{t},a_{t})$, which estimates the expected cumulative reward.
@@ 4 Methodology
We assume that the trajectory data used to train task-specific policy is accessible, and reusing this data allows us to perform zero-shot personalisation i.e., without collecting new environment interactions. A subset of this data is sampled with personalised human feedback scores with more desirable trajectories assigned higher scores. We then infer the human intent (intent-specific policy) using RUDDER (Section 3.2) and dynamically fuse it with the task-specific policy. The overview of our method is illustrated in Fig 1. The subsequent sections provide in-depth details of the components of our personalisation approach.
@@ 4.1 Learning Human Intent using LSTM
At each time step, the state-action vector is fed into the LSTM units. In the case of image inputs, we pre-train a Variational Autoencoder (VAE) to reduce the dimensionality of the state. We use Feature-wise Linear Modulation (FiLM) [18] to modulate the state vector with the action vector. In [18], a FiLM network conditions the feature map of the neural network depending on another input signal. Here, we condition the state feature with the action vector and train the LSTM using the human feedback data as training labels. With this training setup, (More details in supplementary Section B and C) the model approximates the Q-value, which is then converted to the intent-specific policy as described in the next section.
@@ 4.2 Policy Construction
We use a Deep Q Network (DQN) parameterised by $\phi$ to learn the task and it produces the Q-values ($Q$) for each state-action pair. Similarly, the LSTM parameterised by $\psi$ also produces the Q-values ($Q^{\prime}$) corresponding to the human intent. We choose the Boltzmann distribution to convert the Q-values into a policy. Therefore $\pi_{\phi}$ is the task-specific policy and $\pi_{\psi}$ is the intent-specific policy. This choice is motivated by the fact that if the intersection of the support of two policies is an empty set, their product would result in a random policy. By contrast, the Boltzmann distribution assigns at least a small probability to all actions, ensuring that the support of the distribution covers the entire action space.
@@ 4.3 Policy Fusion
Policy fusion is a process of combining two or more policies to produce a new policy.  [20] discussed several approaches through which policy fusion can be performed. We also aim to obtain a personalised policy by fusing the intent-specific and task-specific policies. However, policy fusion in the context of personalisation should satisfy two constraints. Firstly, fused policy should be identical to the task-specific policy if the task-specific and intent-specific policies are identical to each other - a property that we refer to as the invariability constraint. This constraint is crucial because it ensures that when the objectives of the task and the human are aligned, the personalised policy remains consistent and does not deviate during the fusion of identical constituent policies. Correspondingly, we define invariant policies as follows:
@@
Two policies $\pi_{1}(a|s)$ and $\pi_{2}(a|s)$ are said to be invariant policies if
$KL\!\left(\pi_{1}(a|s)\,\middle\|\,\pi_{2}(a|s)\right)=0\quad\forall\,s\in\mathcal{S}.$
@@
By assumption, $\pi_{\phi}(a|s)=\pi_{\psi}(a|s)$. Hence
$KL\!\left(\pi_{\phi}(a|s)\,\middle\|\,\pi_{f}(a|s)\right)=KL\!\left(\pi_{\phi}(a|s)\,\middle\|\,\pi_{\phi}(a|s)\right)=0.$
Therefore, $\pi_{\phi}(a|s)$ and $\pi_{f}(a|s)$ are invariant policies. ∎
@@
Please see Section  E
The above theorem establishes the boundedness of the divergence between the task-specific policy and the personalised policy. Nevertheless, we can consider other choices for policy fusion as studied in the previous work [20], such as the weighted average or the product of the two policies. However, the former does not operate on the common support, and the latter does not satisfy the invariability constraint, even when the policies are exactly the same, as evidenced in the Lemma 2. For a detailed analysis, we refer the reader to the supplementary Section  F.
@@
Let $\pi_{\psi}$ and $\pi_{\phi}$ be intent-specific and task-specific policies respectively, and assume $\pi_{\psi}$ is not a random policy. Then $\pi_{\phi}$ and $\tfrac{1}{Z}\pi_{\psi}\pi_{\phi}$, where $Z$ is a normalising factor, are not invariant policies under any condition.
@@
Please see Section  E
@@ 4.4 Pitfalls of Static Fusion
1:while not donedone do
2:  $Q_{t}\leftarrow[]$ $\triangleright$ DQN Q-values reset
3:  $Q^{\prime}_{t}\leftarrow[]$ $\triangleright$ LSTM Q-values reset
4:  $r_{t}\leftarrow[]$ $\triangleright$ Human-induced reward reset
5:  $Q_{t}\leftarrow\phi(s_{t})$ $\triangleright$ Invoking DQN
6:  $Q^{\prime}_{t}\leftarrow\psi(s_{t},a)$ $\triangleright$ Invoking LSTM
@@
8:  $\pi_{\phi}\leftarrow BoltzmanDistribution(Q_{t},T_{\phi})$
9:  $\pi_{\psi}\leftarrow BoltzmanDistribution(Q^{\prime}_{t},T_{\psi})$
10:  $a_{t}\leftarrow\arg\max_{a\in\mathcal{A}}\sqrt{\pi_{\phi}(a|s_{t})\times\pi_{\psi}(a|s_{t})}$
11:  $r^{\prime}\leftarrow r_{t}-mean(r_{t})$ $\triangleright$ Eq (5)
12:  $g(t)\leftarrow\sum_{t^{\prime}=0}^{t}r^{\prime}(s_{t^{\prime}},a_{t^{\prime}})$
13:  $T_{\psi}\leftarrow max(T_{min},\frac{T_{max}}{1+exp(-m(g-\eta))})$ $\triangleright$ Eq (6)
@@
15:  $t\leftarrow t+1$